OpenStudy (mokeira):

What is the formula for finding: a) Slope of secant line b) Slope of tangent line Please explain in depth

OpenStudy (bradely):

Slope of secant line m=(f(x0)-f(x1))/(x0-x1) Slope of tangent line m=(f(x0+h)-f(x0))/h source: http://www.mathskey.com/question2answer/

OpenStudy (mokeira):

thnks!